OpenStudy (chaise):

Integrate: Mainly interested in the steps involved, as opposed to the answer, so steps involved would be great. Thanks in advance. (1-3x^2)/x dx

OpenStudy (chaise):

\[\int\limits_{-}^{-} \frac{1-3x^2}{x}\]

OpenStudy (anonymous):

use substitution let u=1-3x^2 then du=-6xdx can u do this

OpenStudy (anonymous):

brake it up into two fractions

OpenStudy (anonymous):

\[ \int \frac{1}{x} - \frac{3x^2}{x}dx\]

OpenStudy (anonymous):

you should be able to do it now :)

OpenStudy (chaise):

This helps a lot, thank you. :)

OpenStudy (anonymous):

np.